OpenStudy (anonymous):

1/4z+1/2=3/4 what is z?

OpenStudy (anonymous):

1/4z+1/2=3/4 Let's first multiply everything by 4 to get rid of the pesky denominators z+2=3 now subtract 2 z = 1

OpenStudy (anonymous):

thankyou!